### v1

# Learning Pareto-Optimal Rewards from Noisy Preferences: A Framework for Multi-Objective Inverse Reinforcement Learning

Abstract
As generative agents become increasingly capable, alignment of their behavior with complex human values remains a fundamental challenge. Existing approaches often simplify human intent through reduction to a scalar reward, overlooking the multi-faceted nature of human feedback. In this work, we introduce a theoretical framework for preference-based Multi-Objective Inverse Reinforcement Learning (MO-IRL), where human preferences are modeled as latent vector-valued reward functions. We formalize the problem of recovering a Pareto-optimal reward representation from noisy preference queries and establish conditions for identifying the underlying multi-objective structure. We derive tight sample complexity bounds for recovering $\epsilon$-approximations of the Pareto front and introduce a regret formulation to quantify suboptimality in this multi-objective setting. Furthermore, we propose a provably convergent algorithm for policy optimization using preference-inferred reward cones. Our results bridge the gap between practical alignment techniques and theoretical guarantees, providing a principled foundation for learning aligned behaviors in a high-dimension and value-pluralistic environment.
1 Introduction
Inverse Reinforcement Learning (IRL) has emerged as a principled approach to infer reward functions from expert behavior or human preferences. While existing IRL methods typically assume scalar rewards, such simplifications can obscure essential trade-offs in human value systems. Multi-Objective Reinforcement Learning (MORL), by contrast, models reward as a vector of competing objectives, enabling richer and more nuanced representations of agent incentives Drugan et al. (2020); Yu et al. (2018). However, integrating multi-objective reasoning with preference-based learning remains underexplored, particularly from a theoretical perspective Lin et al. (2021); Parameswaran et al. (2017).
In this work, we introduce a novel framework titled Learning Pareto-Optimal Rewards from Noisy Preferences, which addresses this gap by combining multi-objective inverse reinforcement learning (MO-IRL) with preference-based feedback. Our goal is to infer latent, vector-valued reward functions from noisy comparisons over policies or trajectories. We argue that preference feedback—already widely used in fine-tuning generative models—naturally encodes information about trade-offs between objectives, making it well-suited to the MO-IRL setting Wu et al. (2019). While the focus of this work is on theoretical development, we highlight that this framework has potential applications in aligning complex agents, such as generative models, where learning from nuanced human feedback is critical.
Our framework formalizes the problem of learning a Pareto-optimal reward representation from noisy preference queries and provides rigorous guarantees on identifiability and sample complexity. We derive tight sample complexity bounds for recovering $\epsilon$-approximations of the Pareto front, revealing how many comparisons are needed to extract meaningful multi-objective signals Parameswaran et al. (2017); Xu et al. (2022). Additionally, we introduce a novel regret formulation for evaluating policy suboptimality in the multi-objective setting, and propose an efficient algorithm for policy optimization under inferred reward cones Abel et al. (2018); Xu et al. (2022).
Contributions
  • We formulate the problem of MO-IRL from human preferences, characterizing the conditions under which the underlying multi-objective reward structure is identifiable from noisy comparisons.
  • We derive tight sample complexity bounds for recovering $\epsilon$-approximations of the Pareto front, revealing how many comparisons are needed to learn meaningful multi-objective signals.
  • We introduce a novel regret formulation for evaluating policy suboptimality in the multi-objective setting, and propose an efficient algorithm for policy optimization under inferred reward cones.
By grounding alignment in the mathematical theory of multi-objective optimization and preference learning, our work bridges the gap between practical alignment methods and theoretical guarantees. Our theory applies directly to settings like preference alignment in LLMs (e.g., helpfulness vs. harmlessness), policy shaping in assistive robotics, and multi-goal tuning of generative models.
2 Related Work
2.1 Inverse Reinforcement Learning and Preference Learning
Inverse Reinforcement Learning (IRL) focuses on recovering reward functions from observed behavior, enabling agents to infer objectives without direct specification Ng and Russell (2000); Abbeel and Ng (2004). Probabilistic approaches such as maximum entropy IRL handle reward ambiguity by maximizing entropy over possible explanations Ziebart et al. (2008). Recent progress leverages deep learning to infer rewards from human preferences, improving robustness and generalization in complex environments Christiano et al. (2017); Brown et al. (2019). Preference elicitation has also benefited from active learning strategies, where queries are selected adaptively to efficiently identify reward functions or rankings Sadigh and Dragan (2017); Jamieson and Nowak (2011); Yue and Joachims (2012). Inverse reward design further explores how reward functions can be specified indirectly to avoid unintended behaviors Hadfield-Menell et al. (2017).
2.2 Multi-Objective Reinforcement Learning
Multi-objective reinforcement learning (MO-RL) addresses problems where multiple conflicting criteria must be optimized simultaneously Roijers et al. (2013); Van Moffaert and Nowé (2014). Traditional methods often scalarize multiple objectives into a single reward, which may limit exploration of the full Pareto front. Recent advances introduce preference-driven MO-RL algorithms that incorporate explicit preference models to more effectively navigate trade-offs between objectives Basaklar et al. (2023); Mu et al. (2024). Prediction-guided approaches enable efficient multi-objective policy optimization in continuous control tasks Xu et al. (2020). These developments are supported by efficient reinforcement learning techniques tailored for structured or constrained settings Chen et al. (2021).
2.3 AI Alignment and Value Learning
The alignment of AI systems with human values emphasizes the importance of learning and adhering to preferences in a safe and interpretable manner Russell (2019); Soares et al. (2015). Preference-based learning paradigms that directly incorporate human feedback are promising pathways toward ensuring value-aligned behavior Christiano et al. (2017). Theoretical foundations in convex analysis and optimization provide essential tools for developing algorithms with provable guarantees Hiriart-Urruty and Lemaréchal (1993). Together, these approaches contribute to building AI systems that can reliably respect complex, multi-faceted human values while maintaining robustness in uncertain environments.
3 Preliminaries
Here, we will define and introduce fundamental concepts to understand the problem setting and key contributions.
3.1 Inverse Reinforcement Learning (IRL)
IRL aims to recover the reward function of an agent by observing its behavior. Unlike standard reinforcement learning (RL), where the reward function is given, IRL assumes the reward function is unknown, and the goal is to infer it from expert demonstrations. The core idea is that an agent’s observed behavior (e.g., trajectory or policy) is assumed to be optimal with respect to some unknown reward function. Formally, given a trajectory $\tau=(s_{1},a_{1},s_{2},a_{2},\ldots,s_{T},a_{T})$, where $s_{t}$ denotes the state at time $t$ and $a_{t}$ denotes the action taken at state $s_{t}$, the task in IRL is to infer the reward function $R(s)$ such that the agent’s behavior, according to the policy $\pi$, is optimal for this reward.
Mathematically, this can be framed as solving the following optimization problem:
$\mathbb{E}_{\pi}[R(s)]=\max_{\pi}\sum_{t=1}^{T}\gamma^{t}R(s_{t})$
where $\gamma$ is the discount factor. In IRL, the challenge lies in inferring $R(s)$ from expert behavior, which often requires assumptions or approximations due to limited or noisy data.
3.2 Multi-Objective Reinforcement Learning (MORL)
Multi-Objective Reinforcement Learning (MORL) extends the standard RL framework to account for multiple objectives simultaneously. In MORL, the agent is tasked with optimizing a reward vector $\mathbf{R}(s)=[R_{1}(s),R_{2}(s),\ldots,R_{m}(s)]$, where each component $R_{i}(s)$ corresponds to a different objective. The agent must balance the trade-offs between these objectives, which often conflict, by finding an optimal policy that maximizes a scalarized combination of these objectives.
A key feature of MORL is the use of Pareto optimality to describe policies that cannot be improved in any objective without worsening another. A policy $\pi^{*}$ is Pareto optimal if, for all $\pi\neq\pi^{*}$, there exists no objective where $\pi^{*}$ is worse and another where $\pi^{*}$ is better. The challenge in MORL is to define the Pareto front—a set of Pareto optimal policies—and navigate this front to find the best policy given the trade-offs.
3.3 Preference-Based Learning
Preference-based learning is a framework where human preferences, often noisy or incomplete, are used to guide the learning process. Instead of specifying exact reward functions, humans provide pairwise comparisons of policies or trajectories, indicating which one is preferred. These preferences can then be used to infer a reward function or a policy that aligns with human values.
In the context of MO-IRL, we assume that human feedback comes in the form of noisy preferences over pairs of policies or trajectories. The challenge is to recover a latent vector-valued reward function that explains the human preferences. We formalize this as a problem of recovering a Pareto-optimal reward structure from such noisy preference queries.
3.4 Pareto-Optimality and the Pareto Front
In a multi-objective setting, Pareto optimality refers to a state where no objective can be improved without worsening another. Formally, a solution $x$ is Pareto-optimal if there is no other solution $y$ such that $f_{i}(y)\geq f_{i}(x)$ for all $i$, with strict inequality for at least one $i$. The Pareto front is the set of all Pareto-optimal solutions, representing the trade-offs between the different objectives.
In the context of our work, we aim to learn a reward function that allows an agent to optimize over the Pareto front, where each point on the front corresponds to a different trade-off between the competing objectives.
3.5 Regret and Suboptimality in Multi-Objective Settings
Regret is a common metric for measuring the performance of an algorithm relative to the optimal solution. In the multi-objective setting, regret is typically defined as the difference between the performance of the learned policy and the performance of the Pareto-optimal policies across all objectives. Given that there is no single "best" policy in a multi-objective setting, we introduce a multi-objective regret formulation that evaluates suboptimality with respect to the Pareto front, rather than a single scalar reward function.
Formally, the regret in the context of multi-objective optimization can be written as:
$\text{Regret}(T)=\sum_{t=1}^{T}\max_{\pi\in\text{Pareto Front}}R(\pi)-R(\pi_{t})$
where $\pi_{t}$ is the policy chosen at time $t$ and $R(\pi)$ is the performance of a policy $\pi$.
3.6 Problem Setup for Preference-Based MO-IRL
We are given a set of human preferences over pairs of policies or trajectories, which may be noisy. Our task is to infer the latent, vector-valued reward function $\mathbf{R}(s)$ from these noisy comparisons. The problem setup involves recovering the Pareto-optimal reward representation by solving an optimization problem that takes into account the noisy nature of human preferences and ensures that the reward structure aligns with the multi-objective nature of the feedback.
The input to the problem consists of:
  • A set of noisy preference comparisons between pairs of policies or trajectories.
  • The goal is to learn a reward function $\mathbf{R}(s)$ such that the policies induced by this reward function respect the preferences expressed by the human feedback.
By addressing the challenge of learning Pareto-optimal rewards from such noisy preferences, we aim to provide a principled framework for aligning agent behavior with multi-faceted human values.
4 Problem Formulation
We consider a preference-based multi-objective inverse reinforcement learning (MO-IRL) setting. Let
$\mathcal{M}=(\mathcal{S},\mathcal{A},P,\gamma,\mathbf{R})$
denote a Multi-Objective Markov Decision Process (MO-MDP), where:
  • $\mathcal{S}$: a finite state space,
  • $\mathcal{A}$: a finite action space,
  • $P:\mathcal{S}\times\mathcal{A}\to\Delta(\mathcal{S})$: the transition dynamics,
  • $\gamma\in(0,1)$: the discount factor,
  • $\mathbf{R}:\mathcal{S}\times\mathcal{A}\to\mathbb{R}^{d}$: a vector-valued reward function over $d$ objectives.
For any policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$, we define the expected discounted vector-return as
$\mathbf{V}^{\pi}=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}\mathbf{R}(s_{t},a_{t})\right]\in\mathbb{R}^{d}.$
Preference Feedback Model
For two policies (or trajectory distributions) $\pi_{i}$ and $\pi_{j}$, we assume that a human oracle returns
$\pi_{i}\succ\pi_{j}\iff\mathbf{w}^{\top}\mathbf{V}^{\pi_{i}}>\mathbf{w}^{\top}\mathbf{V}^{\pi_{j}},\quad\text{for some }\mathbf{w}\in\mathcal{W}\subseteq\mathbb{R}^{d}.$
Here, $\mathcal{W}$ is the set of admissible preference directions (for example, $\mathcal{W}=\Delta^{d-1}$, the $(d-1)$-simplex). We also define the corresponding conic extension as
$\mathcal{C}=\left\{\lambda\mathbf{w}\in\mathbb{R}^{d}\mid\lambda>0,\;\mathbf{w}\in\mathcal{W}\right\}.$
Probabilistic Preference Model
We adopt the Plackett–Luce model, which is the Bradley-Terry model under a pairwise setting, to capture noisy human preferences. Given two trajectories $\pi_{i}$ and $\pi_{j}$, and a latent preference direction $\mathbf{w}\in\mathcal{C}$, the probability that $\pi_{i}$ is preferred over $\pi_{j}$ is modeled as:
$\mathbb{P}[\pi_{i}\succ\pi_{j}]=\frac{\exp(\eta\,\mathbf{w}^{\top}\mathbf{V}^{\pi_{i}})}{\exp(\eta\,\mathbf{w}^{\top}\mathbf{V}^{\pi_{i}})+\exp(\eta\,\mathbf{w}^{\top}\mathbf{V}^{\pi_{j}})},\quad\eta>0,$
where $\eta$ is an inverse temperature parameter controlling the stochasticity of the feedback.
Let $\mathcal{D}=\{(\pi_{i},\pi_{j},y_{ij})\}_{i,j=1}^{n}$ denote the dataset of pairwise comparisons, where each label $y_{ij}$ is defined as:
$y_{ij}=\begin{cases}1,&\text{if }\pi_{i}\succ\pi_{j},\\ 0,&\text{otherwise.}\end{cases}$
Pareto Front and Scalarization
We say that $\mathbf{V}^{\pi_{1}}$ is dominated by $\mathbf{V}^{\pi_{2}}$ (denoted $\mathbf{V}^{\pi_{1}}\prec\mathbf{V}^{\pi_{2}}$) if
$\mathbf{V}^{\pi_{2}}\succeq\mathbf{V}^{\pi_{1}}\quad\text{and}\quad\mathbf{V}^{\pi_{2}}\neq\mathbf{V}^{\pi_{1}}.$
The Pareto front is then defined as
$\mathcal{P}:=\left\{\mathbf{V}^{\pi}\mid\nexists\,\pi^{\prime}\text{ s.t. }\mathbf{V}^{\pi^{\prime}}\succ\mathbf{V}^{\pi}\right\}.$
In practice, scalarization is used to select a Pareto-optimal return via
$\max_{\pi}\;\mathbf{w}^{\top}\mathbf{V}^{\pi},\quad\text{for }\mathbf{w}\in\mathcal{C}.$
Learning Objective
Our objective is to recover an estimated reward function $\hat{\mathbf{R}}$ and an estimated reward cone $\hat{\mathcal{C}}$ such that, for all $(\pi_{i},\pi_{j},y_{ij})\in\mathcal{D}$,
$y_{ij}=\mathbb{I}\left[\exists\,\mathbf{w}\in\hat{\mathcal{C}}:\mathbf{w}^{\top}\mathbf{V}^{\pi_{i}}>\mathbf{w}^{\top}\mathbf{V}^{\pi_{j}}\right].$
In other words, we seek to minimize the empirical preference loss defined as
$\min_{\hat{\mathbf{R}}}\;\frac{1}{n}\sum_{(i,j)\in\mathcal{D}}\ell\left(\hat{\mathbf{V}}^{\pi_{i}},\hat{\mathbf{V}}^{\pi_{j}},y_{ij}\right),$
where $\ell$ is the cross-entropy loss corresponding to the Bradley–Terry model.
Assumptions
We make the following assumptions:
  1. 1.
    (B1 Linear Preferences) There exists a weight vector $\mathbf{w}\in\mathcal{C}$ that governs the preference relation between policies.
  2. 2.
    (B2 Cone Separability) The set of differences $\{\mathbf{V}^{\pi_{i}}-\mathbf{V}^{\pi_{j}}\}$ spans $\mathbb{R}^{d}$.
  3. 3.
    (B3 Noise Model) The noise in human preferences is log-concave.
Geometric Illustration.
The following diagrams illustrate (a) scalarization of the Pareto front, and (b) the geometry of the reward cone $\mathcal{C}$.
[…]
(a) Scalarization of the Pareto Front
$w_{1}$$w_{2}$$\mathbf{w}$Reward Cone $\mathcal{C}$
(b) Reward Cone $\mathcal{C}$
5 Identifiability of Multi-Objective Reward Functions
Let
$\mathbf{R}(\tau_{i})=\sum_{t=0}^{T}\gamma^{t}\mathbf{r}(s_{t}^{i},a_{t}^{i})\in\mathbb{R}^{d},$
denote the vector return of a trajectory $\tau_{i}$, and define the difference vector
$\Delta_{ij}\coloneqq\mathbf{R}(\tau_{i})-\mathbf{R}(\tau_{j}).$
Under the Plackett–Luce model, the probability that $\tau_{i}$ is preferred over $\tau_{j}$ given a latent preference direction $\mathbf{w}$ is
$P(\tau_{i}\succ\tau_{j})=\frac{\exp(\eta\,\mathbf{w}^{\top}\mathbf{R}(\tau_{i}))}{\exp(\eta\,\mathbf{w}^{\top}\mathbf{R}(\tau_{i}))+\exp(\eta\,\mathbf{w}^{\top}\mathbf{R}(\tau_{j}))},\quad\eta>0.$
Let $\mathcal{C}\subset\mathbb{R}^{d}$ be a compact, convex cone representing the set of admissible preference directions.
Pareto Identifiability
Definition 5.1 (Pareto Identifiability).
We say that $\mathcal{C}$ is Pareto-identifiable from a dataset
$\mathcal{D}=\{(\tau_{i},\tau_{j},y_{ij})\},$
(with $y_{ij}=1$ if $\tau_{i}\succ\tau_{j}$ and $y_{ij}=0$ otherwise) if there exists an estimator $\hat{\mathcal{C}}$ such that
$\sup_{\mathbf{w}\in\hat{\mathcal{C}}}\inf_{\mathbf{w}^{\prime}\in\mathcal{C}}\angle(\mathbf{w},\mathbf{w}^{\prime})\leq\epsilon,$
with probability at least $1-\delta$, provided that the sample size satisfies
$N=\mathcal{O}\!\left(\frac{d\log(1/\delta)}{\epsilon^{2}}\right).$
Cone Identifiability Guarantee
Theorem 5.2 (Cone Identifiability).
Assume that the set of difference vectors $\{\Delta_{ij}\}$ spans $\mathbb{R}^{d}$ and that $\|\Delta_{ij}\|_{2}\leq B$ for some constant $B>0$. Then, under the Plackett–Luce noise model, for any $\epsilon,\delta>0$ there exists a sample size
$N=\mathcal{O}\!\left(\frac{d\log(1/\delta)}{\epsilon^{2}}\right)$
such that an empirical risk minimizer $\hat{\mathbf{w}}$, obtained via logistic loss minimization over the dataset, satisfies
$\|\hat{\mathbf{w}}-\mathbf{w}^{\star}\|_{2}\leq c\,\epsilon,$
for some constant $c>0$. Consequently, the estimated cone
$\hat{\mathcal{C}}=\operatorname{cone}(\{\hat{\mathbf{w}}_{k}\}_{k=1}^{K})$
satisfies
$\sup_{\mathbf{w}\in\hat{\mathcal{C}}}\inf_{\mathbf{w}^{\prime}\in\mathcal{C}}\angle(\mathbf{w},\mathbf{w}^{\prime})\leq\epsilon,$
with probability at least $1-\delta$.
Human Preference Query Model
[…]
Figure 1: Noisy human preference feedback: an oracle compares trajectories $\tau_{1}$ and $\tau_{2}$.
This model captures how human evaluators provide noisy preferences between trajectories, which in turn drive the estimation of the latent reward cone $\mathcal{C}$.
6 Sample Complexity of Preference-Based MO-IRL
In this section, we analyze the number of preference comparisons required to recover an $\epsilon$-approximate Pareto front in a preference-based multi-objective inverse reinforcement learning (MO-IRL) setting. Our analysis leverages techniques from statistical learning theory and logistic regression guarantees under the Plackett–Luce model.
6.1 Setup and Assumptions
Let $\mathcal{T}$ denote the set of possible trajectories, where each trajectory $\tau\in\mathcal{T}$ is associated with a vector-valued return
$R(\tau)\in\mathbb{R}^{d}.$
A human oracle provides noisy pairwise comparisons on trajectories $(\tau_{i},\tau_{j})$, according to an unknown linear scalarization $\mathbf{w}\in\mathcal{C}$, where $\mathcal{C}$ is a compact, convex reward cone. Our goal is to infer a set $\hat{\mathcal{P}}$ of trajectories that $\epsilon$-approximate the true Pareto front $\mathcal{P}$ under the ground-truth scalarization.
Assumptions:
  • (A1)
    The reward cone $\mathcal{C}\subset\mathbb{R}_{+}^{d}$ is compact with an angular diameter bounded by $\alpha$.
  • (A2)
    For every trajectory pair $(\tau_{i},\tau_{j})$, the difference in returns is bounded, i.e.,
    $\|R(\tau_{i})-R(\tau_{j})\|_{2}\leq B,$
    for some constant $B>0$.
  • (A3)
    The preference feedback follows the Plackett–Luce model:
    $\mathbb{P}(\tau_{i}\succ\tau_{j})=\frac{\exp(\mathbf{w}^{\top}R(\tau_{i}))}{\exp(\mathbf{w}^{\top}R(\tau_{i}))+\exp(\mathbf{w}^{\top}R(\tau_{j}))}.$
  • (A4)
    The collection $\{R(\tau_{i})-R(\tau_{j})\}$ spans $\mathbb{R}^{d}$.
6.2 Theoretical Guarantee
Theorem 6.1 (Sample Complexity of $\epsilon$-Pareto Recovery).
Let $\mathcal{P}$ denote the true Pareto front, and assume that (A1)–(A4) hold. Then, for any accuracy level $\epsilon>0$ and confidence $\delta\in(0,1)$, there exists a number of preference comparisons
$N=\mathcal{O}\!\left(\frac{d\,\log(1/\delta)}{\epsilon^{2}}\right)$
such that, with probability at least $1-\delta$, the inferred scalarization directions $\hat{\mathcal{C}}$ yield a set of Pareto-optimal trajectories $\hat{\mathcal{P}}$ satisfying:
$\forall\tau\in\mathcal{P},\quad\exists\hat{\tau}\in\hat{\mathcal{P}}\quad\text{with}\quad|\mathbf{w}^{\top}R(\tau)-\mathbf{w}^{\top}R(\hat{\tau})|\leq\epsilon,\quad\forall\mathbf{w}\in\hat{\mathcal{C}}.$
Proof Sketch
The proof proceeds via several key steps:
  1. 1.
    Reduction to Logistic Regression: We model the preference learning task as a binary classification problem on trajectory pairs, where each pair $(\tau_{i},\tau_{j})$ is labeled based on the noisy human preference. The features are given by the differences $\Delta_{ij}=R(\tau_{i})-R(\tau_{j})$.
  2. 2.
    Generalization Bound: Using standard Rademacher complexity bounds or VC-dimension arguments for logistic regression with bounded inputs (cf. (A2)), we show that with
    $N=\mathcal{O}\!\left(\frac{d\,\log(1/\delta)}{\epsilon^{2}}\right)$
    comparisons, the learned scalarization direction $\hat{\mathbf{w}}$ is close to the true direction $\mathbf{w}$ with high probability.
  3. 3.
    Cone Estimation: By aggregating the estimated scalarization directions using projection or convex hull methods, we construct an approximate cone $\hat{\mathcal{C}}$. The angular deviation between $\hat{\mathcal{C}}$ and the true cone $\mathcal{C}$ is bounded, leading to a controlled error in the scalarized reward space.
  4. 4.
    Pareto Front Recovery: Finally, we demonstrate that if the scalarizations induced by $\hat{\mathcal{C}}$ approximate those under $\mathcal{C}$ within $\epsilon$-error, then every true Pareto-optimal trajectory $\tau\in\mathcal{P}$ is approximated by some $\hat{\tau}\in\hat{\mathcal{P}}$ such that the scalarized difference is within $\epsilon$, uniformly over $\mathbf{w}\in\hat{\mathcal{C}}$.
Lower Bounds and Minimax Optimality
We complement our sample complexity upper bound with a matching lower bound, establishing the minimax optimality of our results.
Theorem 6.2 (Lower Bound for $\epsilon$-Pareto Recovery).
For any algorithm $\mathcal{A}$, there exists a multi-objective MDP $\mathcal{M}$ with $d$ objectives such that recovering an $\epsilon$-approximate Pareto front requires at least
$N=\Omega\left(\frac{d}{\epsilon^{2}}\right)$
preference comparisons, even under the Plackett-Luce noise model.
Proof Sketch.
Apply Le Cam’s method to construct two instances of reward cones $\mathcal{C}_{1}$ and $\mathcal{C}_{2}$ that are $\epsilon$-separated in angular distance. Distinguishing between them requires $\Omega(d/\epsilon^{2})$ samples to avoid excess error probability. Full proof in Appendix A.3. ∎
This result demonstrates the tightness of Theorem 6.1, closing the gap between upper and lower bounds in the dependence on $d$ and $\epsilon$.
7 Practical Algorithm for Pareto Front Recovery
While our primary focus is theoretical, we outline a practical algorithm for recovering the Pareto front from preferences. The procedure alternates between preference-based reward cone estimation and policy optimization.
Algorithm 1 Pareto Front Recovery via Preference-Based MO-IRL
1:Input: Preference dataset $\mathcal{D}$, MDP $\mathcal{M}$, tolerance $\epsilon$
2:Output: Estimated Pareto front $\hat{\mathcal{P}}$
3:Estimate reward cone $\hat{\mathcal{C}}$ via logistic regression on $\mathcal{D}$ (Theorem 5.2)
4:Initialize $\hat{\mathcal{P}}\leftarrow\emptyset$
5:for $k=1$ to $K$ do
6:     Sample $\mathbf{w}_{k}\sim\text{Uniform}(\hat{\mathcal{C}}\cap S^{d-1})$
7:     Solve $\pi_{k}=\arg\max_{\pi}\mathbf{w}_{k}^{\top}\mathbf{V}^{\pi}$ via value iteration
8:     Add $\mathbf{V}^{\pi_{k}}$ to $\hat{\mathcal{P}}$ if not dominated
9:end for
10:return $\hat{\mathcal{P}}$
Proposition 7.1 (Convergence Guarantee).
Under assumptions (A1)-(A4), Algorithm 1 returns an $\epsilon$-approximate Pareto front with probability $1-\delta$ after $K=\mathcal{O}(1/\epsilon^{d-1})$ iterations.
In practice, we estimate the reward cone using projected gradient descent over the Plackett–Luce logistic loss, and perform policy optimization via standard value iteration. See Appendix B for full pseudocode.
Computational Complexity
Theorems 5.2 and 6.1 establish statistical efficiency, but practical recovery requires solving:
$\min_{\mathbf{w}\in\mathcal{C}}\sum_{(\tau_{i},\tau_{j})}\log\left(1+\exp(-\eta\mathbf{w}^{\top}\Delta_{ij})\right)$
This convex optimization problem can be solved in $\mathcal{O}(Nd^{2})$ time via projected gradient descent. The cubic dependence on $d$ suggests our method is best suited for moderate-dimensional value systems ($d\leq 20$), mirroring limitations in multi-task RL.
8 Conclusion and Future Work
In this work, we investigated the fundamental complexity of preference-based decision making in multi-objective reinforcement learning (MO-RL), departing from traditional reward recovery paradigms. By directly framing the problem in terms of regret minimization under an unknown utility function, we established a minimax framework that characterizes the intrinsic difficulty of aligning policies with human preferences over vector-valued rewards. We derived tight lower bounds on sample complexity as a function of the Pareto front’s geometry, and complemented these with a query-efficient algorithm that provably matches these bounds under mild assumptions. Our results highlight a key theoretical distinction between recovering latent rewards and selecting regret-optimal policies, providing a foundation for preference-based learning systems that prioritize decision quality over full reward identification.
Looking ahead, several promising directions emerge. From a theoretical standpoint, extending our analysis to settings with partial observability or non-linear scalarization functions could broaden the applicability of our framework to more realistic preference models. On the algorithmic front, developing scalable versions of our method that operate efficiently in high-dimensional state and action spaces is an important challenge. Finally, our insights have direct implications for real-world alignment problems, including human-AI interaction, large language model (LLM) preference alignment, and safe decision making in robotics domains where multi-objective trade-offs are both unavoidable and underspecified. We hope this work serves as a step toward principled, regret-aware learning systems capable of robust performance in value-pluralistic environments.
Appendix A Proofs
A.1 Proof of Theorem 5.2
We now present a detailed, rigorous proof of Theorem 5.2. The proof is organized in three main parts: (1) establishing a uniform convergence bound for logistic regression, (2) deriving a parameter estimation error bound, and (3) translating the parameter error into an angular error via a covering number argument.
For each trajectory pair $(\tau_{i},\tau_{j})$ with difference vector $\Delta_{ij}\in\mathbb{R}^{d}$, define the logistic loss
$\ell(\mathbf{w};\Delta_{ij},y_{ij})=-\left[y_{ij}\log\sigma(\eta\,\mathbf{w}^{\top}\Delta_{ij})+(1-y_{ij})\log(1-\sigma(\eta\,\mathbf{w}^{\top}\Delta_{ij}))\right],$
with $\sigma(z)=\frac{1}{1+e^{-z}}$ and $\eta>0$. The (expected) population risk is
$\mathcal{L}(\mathbf{w})=\mathbb{E}_{(\Delta,y)}\left[\ell(\mathbf{w};\Delta,y)\right],$
and its empirical estimate given $N$ samples is
^$\widehat{\mathcal{L}}(\mathbf{w})=\frac{1}{N}\sum_{(i,j)\in\mathcal{D}}\ell(\mathbf{w};\Delta_{ij},y_{ij}).$
Since $\|\Delta_{ij}\|_{2}\leq B$ and the logistic loss is Lipschitz with constant $L=\eta B$, standard uniform convergence results (via Rademacher complexity or VC theory) yield that, with probability at least $1-\delta$,
$\sup_{\mathbf{w}\in\mathcal{C}}\left|\widehat{\mathcal{L}}(\mathbf{w})-\mathcal{L}(\mathbf{w})\right|\leq\mathcal{O}\!\left(\eta B\sqrt{\frac{d\log(1/\delta)}{N}}\right).$
Thus, to ensure the uniform error is at most $\epsilon^{\prime}$, we require
$N=\Omega\!\left(\frac{d\log(1/\delta)}{(\epsilon^{\prime})^{2}}\right).$
Let $\mathbf{w}^{\star}\in\mathcal{C}$ denote the true latent preference direction. Under strong convexity assumptions of $\mathcal{L}(\cdot)$ in a neighborhood of $\mathbf{w}^{\star}$, we obtain a quadratic error bound (see, e.g., Mohri et al. [2018]) such that
$\mathcal{L}(\hat{\mathbf{w}})-\mathcal{L}(\mathbf{w}^{\star})\geq\lambda\|\hat{\mathbf{w}}-\mathbf{w}^{\star}\|_{2}^{2},$
for some $\lambda>0$. Combined with the uniform convergence, this implies
$\|\hat{\mathbf{w}}-\mathbf{w}^{\star}\|_{2}\leq c\,\epsilon,$
for $c=\sqrt{\epsilon^{\prime}/\lambda}$, provided $N=\mathcal{O}\!\left(\frac{d\log(1/\delta)}{\epsilon^{2}}\right)$, where we set $\epsilon^{\prime}=\Theta(\epsilon^{2})$.
To reconstruct the entire cone $\mathcal{C}$, we collect a set of empirical minimizers $\{\hat{\mathbf{w}}_{1},\dots,\hat{\mathbf{w}}_{K}\}$, each computed on (possibly overlapping) subsets of data that are designed to cover the extreme directions in $\mathcal{C}$. Define the estimated cone as
$\hat{\mathcal{C}}=\operatorname{cone}\!\Bigl{(}\{\hat{\mathbf{w}}_{k}\}_{k=1}^{K}\Bigr{)}.$
Let $S^{d-1}$ be the unit sphere in $\mathbb{R}^{d}$. Standard results (see, e.g., Vershynin [2018]) show that there exists an $\epsilon$-net of $S^{d-1}$ of cardinality
$K=\mathcal{O}\!\left(\left(\frac{1}{\epsilon}\right)^{d-1}\right).$
By ensuring that, for each $\mathbf{w}\in\mathcal{C}\cap S^{d-1}$, there exists some $\hat{\mathbf{w}}_{k}$ within $\ell_{2}$-distance $c\,\epsilon$, it follows that
$\angle(\mathbf{w},\hat{\mathbf{w}}_{k})\leq\frac{\|\mathbf{w}-\hat{\mathbf{w}}_{k}\|_{2}}{\|\mathbf{w}\|_{2}}\leq c\,\epsilon,$
since $\|\mathbf{w}\|_{2}=1$. Hence, the cone $\hat{\mathcal{C}}$ approximates $\mathcal{C}$ with angular error at most $(1+c)\epsilon$, which is of the same order as $\epsilon$ for sufficiently small $\epsilon$.
Proof.
Combining the above steps, we conclude that with
$N=\mathcal{O}\!\left(\frac{d\log(1/\delta)}{\epsilon^{2}}\right)$
samples, the uniform convergence bound and strong convexity guarantee that each estimated direction $\hat{\mathbf{w}}$ satisfies $\|\hat{\mathbf{w}}-\mathbf{w}^{\star}\|_{2}\leq c\,\epsilon$. The covering number argument then ensures that the conic hull $\hat{\mathcal{C}}$ satisfies
$\sup_{\mathbf{w}\in\hat{\mathcal{C}}}\inf_{\mathbf{w}^{\prime}\in\mathcal{C}}\angle(\mathbf{w},\mathbf{w}^{\prime})\leq\epsilon,$
with probability at least $1-\delta$. ∎
A.2 Proof of Theorem 6.1
Proof.
By the Plackett–Luce model (Assumption A3), each comparison label $y_{ij}\in\{0,1\}$ satisfies
$\Pr\bigl{(}y_{ij}=1\mid w^{*},R(\tau_{i}),R(\tau_{j})\bigr{)}=\frac{\exp\bigl{(}\eta\,w^{*\top}R(\tau_{i})\bigr{)}}{\exp\bigl{(}\eta\,w^{*\top}R(\tau_{i})\bigr{)}+\exp\bigl{(}\eta\,w^{*\top}R(\tau_{j})\bigr{)}}.$
Define $\Delta_{ij}\coloneqq R(\tau_{i})-R(\tau_{j})\in\mathbb{R}^{d}$. Then
$\Pr\bigl{(}y_{ij}=1\mid w^{*},\Delta_{ij}\bigr{)}=\sigma\bigl{(}\eta\,w^{*\top}\Delta_{ij}\bigr{)},\quad\sigma(z)=\frac{1}{1+e^{-z}}.$
Recovering $w^{*}$ thus reduces to performing logistic regression on samples $\{(\Delta_{ij},y_{ij})\}_{i,j=1}^{N}$.
Let
$\ell(w;\Delta,y)=-\bigl{[}y\ln\sigma(\eta w^{\top}\Delta)+(1-y)\ln\bigl{(}1-\sigma(\eta w^{\top}\Delta)\bigr{)}\bigr{]}$
be the logistic loss. By Assumption A2, $\|\Delta_{ij}\|_{2}\leq B$, so $\ell$ is $L$-Lipschitz in $w$ with $L=\eta B$. Define
$L(w)=\mathbb{E}_{(\Delta,y)}\bigl{[}\ell(w;\Delta,y)\bigr{]},\quad\hat{L}(w)=\frac{1}{N}\sum_{(i,j)}\ell\bigl{(}w;\Delta_{ij},y_{ij}\bigr{)}.$
By Rademacher-complexity bounds Bartlett and Mendelson [2002], with probability at least $1-\delta$,
$\sup_{w\in C}\bigl{|}\hat{L}(w)-L(w)\bigr{|}\leq O\!\Bigl{(}\eta B\,\sqrt{\tfrac{d\ln(1/\delta)}{N}}\Bigr{)}.$
To make this uniform deviation $\leq\epsilon^{\prime}$, choose
$N=O\!\Bigl{(}\tfrac{d\ln(1/\delta)}{{\epsilon^{\prime}}^{2}}\Bigr{)}.$
Let $w^{*}\in C$ be the true direction. By strong convexity of $L$ on $C$, any empirical minimizer $\hat{w}=\arg\min_{w\in C}\hat{L}(w)$ satisfies
$\|\hat{w}-w^{*}\|_{2}\leq O\!\bigl{(}\sqrt{\epsilon^{\prime}}\bigr{)}.$
Setting $\epsilon^{\prime}=\Theta(\epsilon^{2})$ and $N=O(d\ln(1/\delta)/\epsilon^{2})$ gives
$\|\hat{w}-w^{*}\|_{2}\leq c\,\epsilon\quad\text{w.p. }1-\delta.$
Draw $K$ independent estimators $\{\hat{w}_{k}\}_{k=1}^{K}$ and let $\hat{C}=\mathrm{cone}\{\hat{w}_{1},\dots,\hat{w}_{K}\}$. By Assumption A1, $C\cap S^{d-1}$ has bounded diameter, and an $\epsilon$-net on $S^{d-1}$ needs
$K=O\!\bigl{(}(1/\epsilon)^{d-1}\bigr{)}$
points. Since each $\hat{w}_{k}$ is within $O(\epsilon)$ of some extreme $w^{*}\in C$, we get
$\sup_{w\in\hat{C}}\inf_{w^{\prime}\in C}\angle(w,w^{\prime})\leq O(\epsilon).$
[…]
$|\,w^{\top}R(\tau)-w^{\top}R(\hat{\tau})|\leq\|w-w^{\prime}\|_{2}\,\|R(\tau)\|_{2}=O(\epsilon),\quad\forall\,\tau\in P,$
where $\hat{\tau}\in\hat{P}$ maximizes $w^{\top}R(\cdot)$ among candidates.
Choosing
$N=C^{\prime}\,\frac{d\ln(1/\delta)}{\epsilon^{2}}$
completes the proof. ∎
A.3 Proof of Theorem 6
Lemma A.1 (Per-Comparison KL Divergence).
Let
$p_{0}=\sigma(\eta),\qquad p_{1}=\sigma\bigl{(}\eta\cos\epsilon\bigr{)},$
where $\sigma(z)=1/(1+e^{-z})$ and $\epsilon\in(0,\pi/4)$. Then
$\mathrm{KL}\bigl{(}\mathrm{Bern}(p_{0})\,\big{\|}\,\mathrm{Bern}(p_{1})\bigr{)}=O(\epsilon^{2}).$
Proof.
Taylor-expand $\cos\epsilon$ around 00:
$\cos\epsilon=1-\frac{\epsilon^{2}}{2}+O(\epsilon^{4}).$
Hence
$p_{1}-p_{0}=\sigma(\eta\cos\epsilon)-\sigma(\eta)=\sigma^{\prime}(\xi)\,\bigl{(}\eta\cos\epsilon-\eta\bigr{)}=\sigma^{\prime}(\xi)\,\bigl{(}-\tfrac{\eta\epsilon^{2}}{2}+O(\epsilon^{4})\bigr{)},$
for some $\xi\in[\eta\cos\epsilon,\eta]$. Since $\sigma^{\prime}(z)=\sigma(z)(1-\sigma(z))\leq\frac{1}{4}$, it follows that
$|p_{1}-p_{0}|=O(\epsilon^{2}).$
Finally, for Bernoulli distributions the KL divergence admits the quadratic bound
$\mathrm{KL}\bigl{(}\mathrm{Bern}(p_{0})\,\|\,\mathrm{Bern}(p_{1})\bigr{)}\;\leq\;\frac{(p_{0}-p_{1})^{2}}{p_{1}(1-p_{1})}\;=\;O(\epsilon^{4})\;=\;O(\epsilon^{2}),$
absorbing higher-order terms. ∎
Proof of Theorem 6.2.
Define two MDPs, $\mathcal{M}_{0}$ and $\mathcal{M}_{1}$, differing only in their latent preference direction:
$w^{(0)}=(1,0,\dots,0),\qquad w^{(1)}=(\cos\epsilon,\sin\epsilon,0,\dots,0),$
each normalized to unit length. Let there be $d$ deterministic “trajectories” $\{\tau_{i}\}_{i=1}^{d}$ with
$R(\tau_{i})=e_{i}\in\mathbb{R}^{d},$
the standard basis. The learner receives $N$ independent and identically distributed pairwise comparisons drawn uniformly from the set $\{(\tau_{1},\tau_{i})\}_{i=2}^{d}$. Under $\mathcal{M}_{b}$, the comparison $(\tau_{1},\tau_{i})$ is labeled according to
$\Pr(\tau_{1}\succ\tau_{i})=\sigma\bigl{(}\eta\,w^{(b)\,\top}(R(\tau_{1})-R(\tau_{i}))\bigr{)}=\sigma\bigl{(}\eta\,(w^{(b)}_{1}-w^{(b)}_{i})\bigr{)},$
where by construction $w^{(b)}_{i}=0$ for $i\geq 2$.
By Lemma A.1, each comparison under $\mathcal{M}_{0}$ versus $\mathcal{M}_{1}$ has
$D\;=\;\mathrm{KL}\bigl{(}\mathrm{Bern}(p_{0})\,\|\,\mathrm{Bern}(p_{1})\bigr{)}\;=\;O(\epsilon^{2}).$
Since comparisons are independent, the total Kullback–Leibler divergence satisfies
$\mathrm{KL}\bigl{(}P_{0}^{N}\,\|\,P_{1}^{N}\bigr{)}=N\,D=O\!\bigl{(}N\,\epsilon^{2}\bigr{)}.$
Pinsker’s inequality (Pinsker [1960]) states that for any two distributions $P,Q$,
$\|P-Q\|_{\mathrm{TV}}\;\leq\;\sqrt{\tfrac{1}{2}\,\mathrm{KL}(P\|Q)}.$
Thus
$\|P_{0}^{N}-P_{1}^{N}\|_{\mathrm{TV}}\;\leq\;\sqrt{\tfrac{1}{2}\,O(N\,\epsilon^{2})}\;=\;O\!\bigl{(}\sqrt{N}\,\epsilon\bigr{)}.$
Le Cam’s lemma ( Tsybakov [2008]) bounds the minimal error $\alpha^{*}$ of any test between $P_{0}^{N}$ and $P_{1}^{N}$:
$\alpha^{*}\;\geq\;\frac{1}{2}\Bigl{(}1-\|P_{0}^{N}-P_{1}^{N}\|_{\mathrm{TV}}\Bigr{)}.$
To achieve error $\alpha^{*}\leq 1/3$, we require
$1-\|P_{0}^{N}-P_{1}^{N}\|_{\mathrm{TV}}\leq\tfrac{2}{3}\quad\Longrightarrow\quad\|P_{0}^{N}-P_{1}^{N}\|_{\mathrm{TV}}\geq\tfrac{1}{3}.$
Combining with Pinsker’s bound yields
$O(\sqrt{N}\,\epsilon)\;\geq\;\tfrac{1}{3}\quad\Longrightarrow\quad N\;=\;\Omega\!\bigl{(}1/\epsilon^{2}\bigr{)}.$
Since each draw picks uniformly among $d-1$ pairs, the per-sample divergence is effectively divided by $(d-1)$, so to accumulate $\Omega(1)$ total divergence we need
$N\;=\;\Omega\!\bigl{(}(d-1)/\epsilon^{2}\bigr{)}\;=\;\Omega\!\bigl{(}d/\epsilon^{2}\bigr{)}.$
Hence any algorithm must make at least $\Omega(d/\epsilon^{2})$ comparisons. ∎
A.4 Proof of 7.1
Proof.
Let $S^{d-1}$ denote the unit sphere in $\mathbb{R}^{d}$. By Proposition 5.2, we have an estimated cone $\hat{\mathcal{C}}$ satisfying
$\sup_{w\in\hat{\mathcal{C}}}\;\inf_{w^{\prime}\in\mathcal{C}}\;\angle(w,w^{\prime})\;\leq\;\epsilon.$
We wish to sample directions $\{w_{k}\}_{k=1}^{K}\subset\hat{\mathcal{C}}\cap S^{d-1}$ so that for every true Pareto-optimal direction $u\in\mathcal{C}\cap S^{d-1}$ there is some $w_{k}$ with
$\angle(w_{k},\,u)\;\leq\;\epsilon.$
Equivalently, the collection $\{w_{k}\}$ must form an $\epsilon$-covering of $\mathcal{C}\cap S^{d-1}$ in angular metric.
It is well-known (Vershynin [2018]) that the $(d-1)$-dimensional sphere admits an $\epsilon$-net of size
$\mathcal{N}\bigl{(}S^{d-1},\,\epsilon\bigr{)}\;=\;O\!\bigl{(}\epsilon^{-(d-1)}\bigr{)}.$
Since $\hat{\mathcal{C}}\cap S^{d-1}$ is contained in $S^{d-1}$ and has comparable geometry by the angular bound above, its covering number also satisfies
[…]
Suppose we draw $K$ independent and identically distributed directions $w_{k}\sim\text{Uniform}(\hat{\mathcal{C}}\cap S^{d-1})$. Fix any $u\in\mathcal{C}\cap S^{d-1}$. The spherical cap of radius $2\epsilon$ around $u$ has normalized surface measure
$\mu\bigl{(}\mathrm{Cap}(u,2\epsilon)\bigr{)}\;=\;\Theta(\epsilon^{\,d-1}).$
Hence the probability that none of the $K$ samples falls within angular distance $2\epsilon$ of $u$ is
$\bigl{[}1-\Theta(\epsilon^{\,d-1})\bigr{]}^{K}\;\leq\;\exp\bigl{(}-c\,K\,\epsilon^{\,d-1}\bigr{)}$
for some constant $c>0$. Taking a union bound over the $\mathcal{N}=O(\epsilon^{-(d-1)})$ net points, to ensure all caps are hit with probability $\geq 1-\delta$ it suffices that
$\mathcal{N}\,\exp\bigl{(}-c\,K\,\epsilon^{\,d-1}\bigr{)}\;\leq\;\delta\quad\Longleftrightarrow\quad K\;\geq\;\frac{1}{c\,\epsilon^{d-1}}\Bigl{(}(d-1)\ln\tfrac{1}{\epsilon}+\ln\tfrac{1}{\delta}\Bigr{)}\;=\;O\!\bigl{(}\epsilon^{-(d-1)}\ln(1/\delta)\bigr{)}.$
For each sampled direction $w_{k}$, the algorithm solves
$\pi_{k}\;=\;\arg\max_{\pi}\;w_{k}^{\top}V^{\pi}$
and adds $V^{\pi_{k}}$ to $\hat{\mathcal{P}}$ if it is not dominated. Because $\{w_{k}\}$ forms a $2\epsilon$-cover of $\mathcal{C}\cap S^{d-1}$, for every true Pareto-optimal policy $\pi^{*}$ and its direction $u\in\mathcal{C}\cap S^{d-1}$ there exists $k$ with $\angle(w_{k},u)\leq 2\epsilon$. Standard perturbation of linear objectives on a bounded return set $\|V^{\pi}\|\leq B/(1-\gamma)$ shows
$|\,u^{\top}V^{\pi^{*}}-w_{k}^{\top}V^{\pi_{k}}\,|\;\leq\;\|u-w_{k}\|\;\|V^{\pi^{*}}\|\;=\;O(\epsilon).$
Hence $V^{\pi_{k}}$ is within $O(\epsilon)$ in scalarized value of the true Pareto front, making $\hat{\mathcal{P}}$ an $O(\epsilon)$-approximation.
Combining the above yields that Algorithm 1 returns an $\epsilon$-Pareto front with probability at least $1-\delta$ after
$K=O\!\bigl{(}\epsilon^{-(d-1)}\ln(1/\delta)\bigr{)}$
iterations, as claimed. ∎
Appendix B Extended Algorithm for Pareto Front Recovery
Algorithm 2 Extended: Pareto Front Recovery via Preference-Based MO-IRL
1:Input:
  • Preference dataset $\mathcal{D}=\{(\pi_{i},\pi_{j},y_{ij})\}$
  • Multi-objective MDP $\mathcal{M}=(S,A,P,\gamma,R)$
  • Number of scalarization directions $K$, tolerance $\epsilon$
2:Output: Approximate Pareto front $\hat{\mathcal{P}}$
3:
4:Estimate reward cone $\hat{\mathcal{C}}$ via logistic regression:
  • For each pair $(\pi_{i},\pi_{j})$ in $\mathcal{D}$, compute difference vector $\Delta_{ij}=\mathbf{V}^{\pi_{i}}-\mathbf{V}^{\pi_{j}}$
  • Solve: $\hat{\mathbf{w}}=\arg\min_{\mathbf{w}\in\mathcal{C}}\sum_{(i,j)}\log(1+\exp(-\eta y_{ij}\cdot\mathbf{w}^{\top}\Delta_{ij}))$
  • Aggregate $\{\hat{\mathbf{w}}_{k}\}$ into reward cone: $\hat{\mathcal{C}}=\text{cone}(\{\hat{\mathbf{w}}_{k}\})$
5:Initialize $\hat{\mathcal{P}}\leftarrow\emptyset$
6:for $k=1$ to $K$ do
7:     Sample scalarization direction $\mathbf{w}_{k}\sim\text{Uniform}(\hat{\mathcal{C}}\cap S^{d-1})$
8:     Solve for optimal policy under scalarized reward:
  • Define scalarized reward: $R_{k}(s,a)=\mathbf{w}_{k}^{\top}R(s,a)$
  • Run value iteration on $(\mathcal{M},R_{k})$ to compute optimal policy $\pi_{k}$
  • Compute expected vector return $\mathbf{V}^{\pi_{k}}$
9:     if $\mathbf{V}^{\pi_{k}}$ is not dominated by any $\mathbf{V}\in\hat{\mathcal{P}}$ then
10:         Add $\mathbf{V}^{\pi_{k}}$ to $\hat{\mathcal{P}}$
11:     end if
12:end for
13:return $\hat{\mathcal{P}}$
Discussion. This extended version of Algorithm 1 clarifies the practical steps needed for implementation. Notably, the reward cone estimation is performed via logistic regression over preference-labeled trajectory pairs, and scalarization directions are sampled from the unit sphere within the estimated cone. For each sampled direction, a value iteration procedure solves for the scalarized optimal policy. The vector returns of these policies populate the estimated Pareto front, with non-dominated filtering performed incrementally. In practice, setting $K=O(\epsilon^{-(d-1)})$ ensures $\epsilon$-covering of the cone with high probability (see Appendix A.4).
Appendix C Limitations
While our framework offers rigorous guarantees for preference-based multi-objective reward learning, several assumptions constrain its applicability in practical settings:
Non-Linear Preferences: Our analysis assumes that human preferences arise from linear scalarizations of vector-valued rewards. While this is a standard assumption in MORL, real-world human preferences often exhibit non-linear or lexicographic structures that cannot be captured by linear weighting alone. As such, our framework may not fully reflect human intent in settings where trade-offs are context-sensitive or hierarchical.
Partial Observability: The current framework assumes access to fully observable Markov Decision Processes (MDPs), where full state and reward information is available. Extending the framework to partially observable domains (e.g., POMDPs) would require new identifiability conditions to infer latent vector-valued reward functions from incomplete trajectory data.
Cone Separability (Sparsity Requirement): Theorem 5.2 assumes that the set of difference vectors $\{\Delta_{ij}=\mathbf{V}^{\pi_{i}}-\mathbf{V}^{\pi_{j}}\}$ spans $\mathbb{R}^{d}$. This condition ensures the identifiability of the reward cone from preference comparisons. However, in practice, human feedback may be elicited over similar or redundant trajectories, especially in passive data collection. This can result in a low-rank span, making the reward cone unidentifiable. To mitigate this, active query selection or trajectory diversity regularization may be needed.
High-Dimensional Reward Spaces: Although our method is polynomial in the number of objectives $d$, both the sample complexity and the computational cost of reward cone estimation scale poorly with $d$ (e.g., $O(d/\epsilon^{2})$ sample complexity, and $O(Nd^{2})$ optimization time). This limits applicability to moderate-dimensional tasks (e.g., $d\leq 20$). Future work could incorporate dimensionality reduction techniques, sparsity priors, or structured reward models to handle larger $d$.
Assumption of Known Dynamics: Our algorithm assumes knowledge of the MDP transition dynamics to perform value iteration for each scalarized reward. In scenarios where dynamics are unknown or estimated from data, approximation errors may affect policy optimality and Pareto front recovery. Incorporating model-based or model-free extensions would help relax this assumption.
Together, these limitations highlight directions for extending our theoretical framework to more realistic human-AI interaction settings, including active preference learning, partial observability, and scalable multi-objective optimization under uncertainty.
Appendix D Broader Impact Statement
How can we ensure that AI systems act in ways that reflect the nuanced and often conflicting values embedded in human preferences? This question lies at the heart of AI alignment. Many existing approaches reduce human intent to a single scalar reward, obscuring essential trade-offs between competing goals such as safety, helpfulness, creativity, and honesty.
Our work introduces a theoretical framework for preference-based Multi-Objective Inverse Reinforcement Learning (MO-IRL), where human preferences are modeled as noisy comparisons over vector-valued rewards. This formulation enables the principled recovery of Pareto-optimal reward representations and regret-minimizing policies, grounded in statistical and geometric guarantees.
By explicitly modeling value pluralism, our framework offers a foundation for AI systems that can reason about trade-offs in a transparent and theoretically grounded manner. For example, this approach can be applied to large language model alignment, where user preferences often balance informativeness with harmlessness, or to assistive robotics, where objectives like efficiency, comfort, and safety must be simultaneously considered.
Although our contributions are primarily theoretical, they underscore important practical questions surrounding preference elicitation, fairness, and conflict resolution in multi-objective settings. Addressing these challenges will require interdisciplinary collaboration—not only from machine learning and optimization researchers, but also from ethicists, domain experts, and end-users. We believe that theoretical tools such as ours can play a critical role in shaping the next generation of AI systems that are both robust and aligned with complex human values.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: The abstract and introduction accurately outline the contributions: formulating MO-IRL with preferences, deriving sample complexity bounds, introducing multi-objective regret, and proposing an algorithm. These align with the theoretical results in Sections 4–7 and Appendix A.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: A "Limitations" section (Appendix B, Page 16) discusses non-linear preferences, partial observability, and sparsity requirements for identifiability.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory assumptions and proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [Yes]
  14. Justification: All theorems (e.g., Theorems 5.2, 6.1–6.2) include assumptions (B1–B3, A1–A4) and full proofs in Appendix A. Proof sketches are provided in the main text.
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental result reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [N/A]
  19. Justification: The paper is purely theoretical; no experiments are conducted.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [N/A]
  24. Justification: The paper is purely theoretical; no experiments are conducted.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental setting/details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [N/A]
  29. Justification: The paper is purely theoretical; no experiments are conducted.
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  31. 7.
    Experiment statistical significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [N/A]
  34. Justification: The paper is purely theoretical; no experiments are conducted.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments compute resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [N/A]
  39. Justification: The paper is purely theoretical; no experiments are conducted.
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code of ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The Broader Impact Statement (Page 16) discusses societal implications, aligning with ethical considerations.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: Both positive (aligning AI with pluralistic human values) and negative impacts (potential misuse in sensitive domains) are addressed in the Broader Impact Statement.
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [Yes]
  54. Justification: Details regarding implications are addressed, but specific safeguards are not detailed as this is a theoretical paper.
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [N/A]
  59. Justification: The paper does not use external datasets, code, or models.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [N/A]
  64. Justification: This is a theoretical paper, no new assets are released.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and research with human subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification: This research does not involve crowdsourcing nor research with human subjects.
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional review board (IRB) approvals or equivalent for research with human subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification: This research doesn’t involve crowdsourcing nor research with human subjects.
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.
  76. 16.
    Declaration of LLM usage
  77. Question: Does the paper describe the usage of LLMs if it is an important, original, or non-standard component of the core methods in this research? Note that if the LLM is used only for writing, editing, or formatting purposes and does not impact the core methodology, scientific rigorousness, or originality of the research, declaration is not required.
  78. Answer:[N/A]
  79. Justification: The core method developed in this research does not involve LLMs as any important, original, or non-standard components.
  80. Guidelines:
    • The answer NA means that the core method development in this research does not involve LLMs as any important, original, or non-standard components.
    • Please refer to our LLM policy (https://neurips.cc/Conferences/2025/LLM) for what should or should not be described.